Pharoah’s Soft Lemonade is starting to develop a new product for which the cash fixed costs are expected to be $90,000. The projected EBIT is $135,000, and the Accounting DOL is expected to be 2.0. What is the Cash Flow DOL?
Answer #1

Accounting DOL = 1+ (Fixed Cost+ Depreciation & Amortization)/EBIT

                         2 = 1+($ 90,000 + Depreciation & Amortization)/$135,000

                          2-1 = ($ 90,000 + Depreciation & Amortization)/$135,000

                             1 x $ 135,000 = $90,000 + Depreciation & Amortization

Depreciation & Amortization= $ 135,000 - $ 90,000= $ 45,000

Cash flow DOL= 1+ Fixed Cost/ (EBIT+ Depreciation & Amortization)

                            = 1+ $ 90,000/ ($ 135,000 + $45,000)

                             = 1+ $ 90,000/$ 180,000

                             = 1+ 0.5

                             = 1.5 times